# Extended formulations for induced tree and path polytopes of chordal graphs

### v2

@@ 1 Induced tree polytope of chordal graphs
In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis. Then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs. Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
@@
Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
@@
The induced tree extended incidence vectors of a chordal graph form a Hilbert basis.
Proof. We show that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ form a Hilbert basis of $\mathcal{C}_{T}(G)$. As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each one-dimensional face contains integer points. This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$.
We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors. Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, $(x^{*},y^{*})|_{G\setminus v}$ is an integer point of $\mathcal{C}_{T}(G\setminus v)$ by construction. By the induction hypothesis, let $(x^{*},y^{*})|_{G\setminus v}$ be generated by the induced trees $\Gamma$ with integer coefficients $\lambda$. The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$. For fixed $uv\in\delta(v)$, each operation preserves the value of $(x^{\prime},y^{\prime})|_{H}$ for $H=G\setminus uv$, and preserves nonnegativity and integrality of $\lambda$. We define the slackness $\Delta=\sum_{uv\in\delta(v)}\Delta(u,v)$ for $\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}$. We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$.
Let us consider the following subsets of $\Gamma$, illustrated in Figure 1 :
@@
  • $\Gamma^{u}_{v}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$
Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
We define the following operation to be applied on $(\Gamma,\lambda)$, which decreases $\Delta$ by an integer value.
$\mathbf{\Gamma^{u}}$-Reduction: let $T$ be an element of $\Gamma^{u}$. Set $\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$ increase $\lambda_{T\cup\{v\}}$ and decrease $\lambda_{T}$ by $\alpha$.
By construction, for any $T\in\Gamma^{u}$, $T\cup\{v\}$ induces a tree and that $\mathbf{\Gamma^{u}}$-Reduction preserves integrality of $\lambda$. Moreover, $\xi^{T}|_{H}=\xi^{T\cup\{v\}}|_{H}$ implying that $\lambda(\Psi)|_{H}$ is preserved.
$\mathbf{\Gamma^{u}}$-Reduction applies until $\Gamma^{u}$ is empty or $\Delta(u,v)=0$.
Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{u}_{v})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction. This show how $\Delta(u,v)=0$ can be achieved on all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with integer $\lambda$.
Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis. In fact, this statement is also true for $C_{4}$. However, it is straightforward to see that this property is closed by taking induced subgraphs. A natural question would be to characterize the graphs for which this property holds.
@@
The induced tree polytope and cone of a chordal graphs have IDP.
@@
We now prove that $\mathcal{T}(G)$ characterizes induced tree polytopes in chordal graphs.
@@
Proof. $\Rightarrow"$ The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators that belong to (4). Theorem 1.1 concludes.
$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $C$ induce a hole of size at least 4 in $G$, and $v$ belongs to $C$. By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(C)$ and $x_{v}=2$ and 0 on all other components, we obtain a point which belongs to $\mathcal{T}(G)$ and is not included in the hypercube implying that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary. ∎
The set of inclusion-wise maximal cliques of chordal graphs is polynomial, implying that (1)–(3) has compact size. Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
@@ 2 Induced path polytope of chordal graphs
We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions]. We use this result to prove a technical graph theory lemma.
After that, we show that the extended incidence vectors of induced paths of a chordal graph form a Hilbert basis. This allows us to give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated constraints come in a polynomial number.
@@
The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis is constructive; therefore, it is necessary to build induced paths with "good properties," and this is what the following lemma stands for.
@@
Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$. Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths. By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$.
To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate. Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system and show in the next theorem that it is generated by the Hilbert basis formed by the induced paths’ extended incidence vectors.
@@
$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1. At most two green edges belong to a same induced path; every blue edge is not contained in any induced path containing some other colored edge. This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
@@ 2.1 Induced paths of chordal graphs and Hilbert basis
In this section, $G$ denotes a chordal graph, $v$ a simplicial vertex, $u$ a neighbor of $v$ and $H$ the graph obtained by removing $uv$ from $G$. We denote by $\Psi$ a family of vertex subsets of $G$ inducing paths, and consider $\lambda$ to be an integer vector indexed by the induced paths of $G$. The pair $(\Psi,\lambda)$ acts as the pair $(\Gamma,\lambda)$ used in the previous section.
We prove that the extended incidence vectors of induced paths of chordal graphs form a Hilbert basis. Although the proof follows the same inductive scheme as in the case of induced trees, it is substantially more technical. In particular, it relies on a refined decomposition of the family $\Psi$.
We first introduce this decomposition of $\Psi$, then define five operations acting on $(\Psi,\lambda)$. For each operation, we establish a corresponding lemma showing that it preserves the property that all elements of $\Psi$ induce paths of $G$, the nonnegativity of $\lambda$ and the value of $\lambda(\Psi)|_{H}:=\sum_{P\in\Psi}\lambda_{P}\xi^{P}|_{H}$. Moreover, we precise its impact on the coefficients $\lambda$. Finally, these ingredients are combined to complete the proof.
Let us consider the following partition of $\Psi$:
@@
  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
@@
We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions . We denote by $K_{\Psi}$ the elements of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$:
$K_{\Psi}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}$
$K_{\Psi}$ is considered to be changing accordingly to $\Psi$. We further partition $\Psi_{I}^{u}$ into:
  • $\Psi_{K_{\Psi}}^{u}$ containing the elements of $\Psi$ intersecting $K_{\Psi}$
  • $\Psi_{\overline{K}_{\Psi}}^{u}$ containing the others.
We now define some operations to be applied on ($\Psi$$\lambda)$. We say that such an operation is valid if it preserves the integrality and nonnegativity of $\lambda$, if the vertex subsets added to $\Psi$ induce paths, and if it preserves the value of $\lambda(\Psi)|_{H}$. In each reduction, $\alpha$ is chosen as an integer satisfying $0<\alpha\leq\min\{\lambda_{P}:P\in\Psi,\ \lambda_{P}\text{ is decreased}\}$. This ensures that $\lambda$ stays nonnegative.
$\mathbf{\Psi^{u}}$-Reduction : let $P$ be an element of $\Psi^{u}$. Increase $\lambda_{P\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
$\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{K_{\Psi}}^{u}$. $P$ contains the two inclusion-wise maximal distinct subsets $P_{1}$ and $P_{2}$ inducing paths ending in $u$. Let $P_{2}$ be the one intersecting $N(v)\setminus u$. Increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
$\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$ such that a neighbor $u_{1}$ of $u$ in $P$ is not complete to $K_{\Psi}$. Let $P_{1}$ be its inclusion-wise maximal subset inducing a path ending in $u$ and containing $u_{1}$, and let $P^{\prime}$ be the other inclusion-wise maximal subset inducing a path ending in $u$. Let $P_{2}$ be an element of $\Psi_{2}^{u}$ whose element $u_{2}$ in $K_{\Psi}$ is non-adjacent to $u_{1}$. Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ by $\alpha$.
$\mathbf{\Psi_{u}^{v}}$-Reduction : Let $P$ be a path of $\Psi_{u}^{v}$ such that the neighbor $u_{1}$ of $u$ distinct from $v$ is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$. Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$. Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$.
$\mathbf{\lambda_{\{u,v\}}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$. $P$ contains the two distinct inclusion-wise maximal subset $P_{1}$ and $P_{2}$ both inducing paths ending in $u$. Note that $P_{1}\cup\{v\}$ induces a path. Decrease $\lambda_{u,v}$ and increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$.
Note that in each of the above operations, $\alpha$ is integer, implying that they preserve integrality of $\lambda$. We show that under some conditions, the above reductions are valid. It remains to verify that the sets of vertices built induce paths and that the value of $\lambda(\Psi)|_{H}$ is preserved.

### v3

@@ 1 Induced tree polytope of chordal graphs
In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis, and then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs.
1.1 Induced trees of chordal graphs and Hilbert basis
Given a graph $G$, and a vertex $v$ we denote by $\Omega^{v}$ the set of maximal cliques containing $v$ and by $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\Omega^{v}$ (1)
@@ 1.1 Induced trees of chordal graphs and Hilbert basis
For binary points, Inequalities (1) express the fact that no two edges incident to $v$ and contained in a same triangle belong to an induced tree. Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
@@
The induced tree extended incidence vectors of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{T}(G)$.
Proof. We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors. Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, and $(\lambda,\Gamma)$ an integer nonnegative combination of $G$’s induced trees’ extended incidence vectors satisfying the following properties, where $(x^{\prime},y^{\prime})=\sum_{T\in\Gamma}\lambda_{T}\Gamma_{T}$:
  1. (i)
    $(x^{\prime},y^{\prime})\leq(x^{*},y^{*})$
  2. (ii)
    $(x^{\prime},y^{\prime})|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$
  3. (iii)
    $(\lambda,\Gamma)$ minimize the quantity $\begin{array}[]{l}\\ \Delta(\lambda,\Gamma):=\sum_{e\in\delta(v)}\Delta_{e}(\lambda,\Gamma)\\ \text{where }\Delta_{e}(\lambda,\Gamma):=y_{e}^{*}-y_{e}^{\prime}\ \forall e\in\delta(v)\end{array}$ among the combinations satisfying (i)-(ii).
$\Delta_{uv}(\lambda,\Gamma)$ corresponds to a Manhattan distance to $(x^{*},y^{*})$. An integer nonnegative combination $(\lambda,\Gamma)$ satisfying Conditions (i)-(ii) is given by the induction hypothesis applied on the integer point $(x^{*},y^{*})|_{G\setminus v}$ of $\mathcal{C}_{T}(G\setminus v)$ hence, $(\lambda,\Gamma)$ exists.
We first show that $\Delta_{uv}(\lambda,\Gamma)=0\ \forall uv\in E$. Fix $uv\in E$ and $H=G\setminus uv$, let us consider the following disjoints subsets of $\Gamma$:
@@
  • $\Gamma^{v}_{u}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$.
Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$. We refer to Figure 1 for an illustration of the different types of induced trees considered, it may be used as a visual aid to keep track of the definitions, we provide two different corresponding induced trees in different shades of green; these are the two bottom ones.
Suppose by contradiction that $\Delta_{uv}(\lambda,\Gamma)$ is nonzero and that $\Gamma^{u}$ is nonempty for some neighbor $u$ of $v$. Let $P$ be an element of $\Gamma^{u}$ and set $\alpha=\min(\Delta_{uv}(\lambda,\Gamma),\lambda_{P})$. Note that $P\cup\{v\}$ is an induced tree, therefore decreasing $\lambda_{P}$ and increasing $\lambda_{P\cup\{v\}}$ by $\alpha$ yields a nonnegative integer combination of induced paths’ extended incidence vectors of $G$, say $(\lambda^{\prime},\Gamma^{\prime})$. As $\xi^{P}|_{G\setminus uv}=\xi^{P\cup\{v\}}|_{G\setminus uv}$, $(\lambda^{\prime},\Gamma^{\prime})$ satisfies Conditions (i)-(ii) and $\Delta_{e}(\lambda,\Gamma)=\Delta_{e}(\lambda^{\prime},\Gamma^{\prime})\ \forall e\in\delta(v)\setminus\{uv\}$. Moreover, $\Delta_{uv}(\lambda^{\prime},\Gamma^{\prime})=\Delta_{uv}(\lambda,\Gamma)-\alpha$, Condition (iii) applied to $(\lambda,\Gamma)$ implies that $\alpha=0$. Therefore, either $\Gamma^{u}$ is empty or $\Delta_{uv}(\lambda,\Gamma)=0$. Suppose that $\Gamma^{u}$ is empty, we now show that $\Delta_{uv}(\lambda,\Gamma)=0$ also holds.
Given a family $\Gamma^{\prime}$ of induced trees, we denote by $\lambda(\Gamma^{\prime})$ the sum of $\lambda$’s components associated with elements of $\Gamma^{\prime}$.
As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{v}_{u})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
The first equality is given by the fact that the only elements of $\Gamma$ contributing to $y_{e}^{\prime},$ with $e\in\delta(u,N(v)\setminus\{u\})$ are the ones of $\Gamma^{u}_{2}$ moreover, each element of $\Gamma^{u}_{2}$ contributes to at most one $y_{e}$ for $e\in\delta(u,N(v)\setminus u)$. The inequality is given by inequality (1) associated with vertex $u$ and clique $N[v]$ . The following equality holds by Condition (ii). The two last equalities hold by the facts that the elements of $\Gamma$ contributing to $x_{u}^{\prime}$ are the ones in $\Gamma^{u}\cup\Gamma^{u}_{2}\cup\Gamma^{v}_{u}$, but $\Gamma^{u}$ is supposed to be empty, and that the only elements of $\Gamma$ contributing to $y_{uv}^{\prime}$ are in $\Gamma^{v}_{u}$.
This shows that $\Delta_{uv}(\lambda,\Gamma)=0$ holds for all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with nonnegative integer coefficients $\lambda$ on induced tree’s extended incidence vectors.
Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis. In fact, this statement is also true for $C_{4}$. Moreover, it is straightforward to see that this property is closed under taking induced subgraphs. A natural question would be to give a forbidden induced subgraph characterization of the graphs for which this property holds. A minimal forbidden such graph is given by the complement graph of a hole of length 7.
1.2 Induced trees of chordal graphs and integral systems
@@
The induced tree polytope of chordal graphs have IDP.
@@
Proof. $\Rightarrow"$ Suppose $G$ to be chordal. The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators which belong to (4). As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each of its one-dimensional face contains integer points. This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$. Therefore, Theorem 1.1 gives that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ generate $\mathcal{C}_{T}(G)$. Extended incidence vectors of induced trees belong to (4) and hence are extreme points of $\mathcal{T}(G)$. It remains to show that $\mathcal{T}(G)$ is bounded. For each vertex $v$ of $G$, there exists a perfect elimination ordering ending in $v$, say $(v_{1},\dots,v_{n},v)$ (it is a consequence of the fact that minimal separators are cliques [Dirac1961]). By definition of perfect elimination orderings, $N(v_{i})\cap\{v_{i+1},\dots,v_{n},v\}$ is a clique and the following inequalities are valid for $\mathcal{C}_{T}(G)$ :
$y(\delta(v_{i},\{v_{i+1},\dots,v_{n},v\}))\leq x_{v_{i}}\quad\forall i\in\{1,\dots,n\}.$
Therefore, summing the previous family of inequalities with x(V)y(E)=1x(V)-y(E)=1 yields $x_{v}\leq 1$. Moreover, $y$ variables are sandwiched between 0 and $x$ variables.
$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $H$ induce a hole of length at least 4 in $G$, and $v$ belongs to $H$. By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(H)$, $x_{v}=2$ and 0 on all other components, we obtain a point $(x,y)$ which belongs to $\mathcal{T}(G)$ and is not included in the hypercube. This implies that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary. ∎
The set of inclusion-wise maximal cliques of chordal graphs being polynomial (1)–(3) has compact size. Therefore, optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
@@ 2 Induced path polytope of chordal graphs
We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions] and use this result to prove a technical lemma. Our proof of the fact that induced paths’ extended incidence vectors of chordal graphs form Hilbert bases follows a similar scheme to the one of the previous section.
We define $(\lambda,\Psi)$ in a similar way to what we did with $(\lambda,\Gamma)$ in the proof of Theorem 1.1. However, the proof of the analogous result for induced path requires a closer study of $(\lambda,\Psi)$. Therefore, we first exhibit a set of properties of $(\lambda,\Psi)$ in the form of lemmas before diving into the proof itself.
In the second section, we deduce an exponential size description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated constraints come in a polynomial number.
@@
The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis requires to "reorganize" the induced path of a conic combination. We mean by reorganize that we build a new induced path having desirable properties from other induced path. This is what stands for the following lemma.
@@
Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, and such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$. Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path and that $u_{1}$ is not adjacent to $u_{2}$. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either their intersection contains $u$ and another vertex, or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle with vertex set $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the edges in $E(P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\})$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of length 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths. By chordality of $G$ and Lemma 2.1, $u_{2}$ is adjacent to $u_{1}$, a contradiction.
To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate. Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system.
$y\Biggl(\delta\Bigl(w,C(K\cup\{w\})\Bigr)\Biggr)+2y\Biggl(\delta\Bigl(w,K\Bigr)\Biggr)$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
@@
For binary points, Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1. At most two green edges belong to a same induced path; given a blue edge $e$, for any other colorful edge $e^{\prime}$, there exists a triangle containing both, therefore no induced path $P$ verifies $\{e,e^{\prime}\}\subseteq E(P)$. This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
Observation 2.3.
The induced paths’ extended incidence vectors of a chordal graph $G$ belong to $\mathcal{C}_{P}(G)$.
@@ 2.1 Induced paths of chordal graphs and Hilbert basis
This section is dedicated to the proof of the following conjecture.
Conjecture 2.4.
The induced path extended incidence vectors of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{P}(G)$.
We prove Conjecture 2.4 by induction over the vertex set of any chordal graph $G$. Note that the initilization is given by the single vertex graph for which $\mathcal{C}_{P}(G)$ is the positive half line trivially generated by a unit vector. The induction proof consists in considering an integer point $(x^{*},y^{*})$ in $\mathcal{C}_{P}(G)$, a simplicial vertex $v$ of $G$ (which exists by chordality of $G$) and a nonnegative integer combination of $G$’s induced paths’ extended vectors say, the induced path $\Psi$ with coefficient $\lambda$ satisfying the following properties where $(x^{\prime},y^{\prime})=\sum_{P\in\Psi}\lambda_{P}\xi^{P}$:
  1. (i)
    $(x^{\prime},y^{\prime})\leq(x^{*},y^{*})$
  2. (ii)
    $(x^{\prime},y^{\prime})|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$
  3. (iii)
    $(\lambda,\Psi)$ minimize the quantity $\begin{array}[]{l}\\ \Delta(\lambda,\Psi):=\sum_{e\in\delta(v)}\Delta_{e}(\lambda,\Psi)\\ \text{where }\Delta_{e}(\lambda,\Psi):=y_{e}^{*}-y_{e}^{\prime}\ \forall e\in\delta(v)\end{array}$ among the ones that satisfy Conditions (i)-(ii).
Note that $(x^{*},y^{*})|_{G\setminus v}$ belong to $\mathcal{C}_{P}(G\setminus v)$, therefore the induction hypothesis implies that $(x^{*},y^{*})|_{G\setminus v}$ is obtained as an integer conic combination of induced paths of $G\setminus v$. This combination, say $(\Psi^{\prime},\lambda^{\prime})$, satisfies Conditions (i)-(ii) but is somehow the worst in terms of Condition (iii) as $\Delta(\lambda^{\prime},\Psi^{\prime})=y^{*}(\delta(v))$. Nevertheless, this proves that $(\lambda,\Psi)$ exists. For the sake of simplicity, we consider that when $\lambda_{P}=0$, $P$ does not belong to $\Psi$.
We first propose a decomposition of $\Psi$, and in the form of lemmas, we exhibit some properties satisfied by $(\lambda,\Psi)$. Given $u\in N(v)$, let us decompose $\Psi$ into the following subsets:
@@
  • $\Psi_{2}^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and contain another vertex $w$ in $N(v)$
  • $\Psi_{I_{2}}^{u}$ is the subset of $\Psi$ whose element’s interior contain $u$ and another element $w$ of $N(v)$.
  • $\Psi_{I_{1}}^{u}$ is the subset of $\Psi$ whose elements’s interior contain $u$ but no other vertex of $N(v)$
  • $\Psi_{u}^{v}$ is the subset of $\Psi$ whose elements contain both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
@@
We denote by $K_{2}^{u}$ the subset of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$, that is:
$K_{2}^{u}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}.$
As the only induced path of $\Psi$ containing the edge $uv$ belong to $\Psi^{v}_{u}$ the following holds:
$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions.